OpenStudy (anonymous):

Use your knowledge of the derivative to compute the value of the limit given below: \lim_{h \rightarrow 0} \frac{\sqrt[1]{(x + h)^{3}} - \sqrt[1]{x^{3}}}{h} . The derivative is \frac{dy}{dx}= .

OpenStudy (ash2326):

\[ \lim_{h \rightarrow 0} \frac{\sqrt[1]{(x + h)^{3}} - \sqrt[1]{x^{3}}}{h} \] This is the derivative of \[\sqrt {x^3}\] Can you find now??

OpenStudy (dumbcow):

you can also write it as derivative of \[\large x^{3/2}\]

OpenStudy (anonymous):

im trying...not getting the correct answer

OpenStudy (dumbcow):

\[\frac{d}{dx} x^{n} = n*x^{n-1}\]

OpenStudy (ash2326):

\[\frac{d}{dx} x^{3/2}=\frac{3}{2} x^{\frac{3}{2}-1}=\frac{3}{2} x^{\frac{1}{2}} \]

OpenStudy (anonymous):

3/2x^(1/2) not correct :(

OpenStudy (ash2326):

Are you checking on a website?

OpenStudy (anonymous):

my homwork is online

OpenStudy (ash2326):

Try \[\frac{3}{2} \sqrt x\]

OpenStudy (anonymous):

3/2x√x is the same as 3/2*x^(1/2) and my online hw says its wrong

razor99 (razor99):

hey ash